Leo Inc. is buying new equipment that has the following cash flows:

Year

0

1

2

3

Cash Flow

-$600

$350

$220

$110

What is the NPV if the interest rate is $6%?

$18.35

$618.35

-$74.01

$79.33

$30.16
Answer #1

The correct answer is $18.35

Net Present Value = Total of Present Value of Cash inflows - Initial Investment.

Years Calculation of Net Present Value Cash flows Discounting ($) Factor @ 6% Present Value($) 350 0.9434 330.1887 220 0.8900

Cash flow at Year 0 is the initial investement for the equipment. Therefore it is shown by a negative sign as it is an outflow.

Note - How did e calculate the discounting factors @6%.

Year 1 = 1/1.06

= 0.9434

Year 2 = 0.9434 /1.06

= 0.8900

Year 3 = 0.8900 /1.06

= 0.8396